Strassel Investors buys real estate, develops it, and resells it for a profit. A new property is available, and Bud Strassel, the president and owner of Strassel Investors, believes if he purchases and develops this property it can then be sold for $155000. The current property owner has asked for bids and stated that the property will be sold for the highest bid in excess of $100000. Two competitors will be submitting bids for the property. Strassel does not know what the competitors will bid, but he assumes for planning purposes that the amount bid by each competitor will be uniformly distributed between $100000 and $145000.

a. Develop a worksheet that can be used to simulate the bids made by the two competitors. Strassel is considering a bid of $125000 for the property. Using a simulation of 1000 trials, what is the estimate of the probability Strassel will be able to obtain the property using a bid of $125000? Round your answer to 1 decimal place. Enter your answer as a percent.

b. Use the simulation model to compute the profit for each trial of the simulation run. With maximization of profit as Strassel’s objective, use simulation to evaluate Strassel’s bid alternatives of $125000, $135000, or $145000. What is the recommended bid, and what is the expected profit?
Answer #1

A.

The total interval is 145000-100000 = 45000

The other 2 winners bid uniformly between the start and end points, so probabilities are uniformly distributed for win/loss

If Strassel bid is 125000

therefore Strassel wins (125000-100000)/45000 = 55.6% times against 1 bidder

Against both bidders, Strassel would have to win against each of them independently, therefore win 55.6% * 55.6% = 30.9% of the times

If Strassel bid is 135000

Against both bidders, using same logic as above, Strassel would win 77.8% * 77.8% = 60.5% of the times

If Strassel bid is 145000

Since Strassel has bid the highest, against both bidders, using same logic as above, Strassel would win 100% of the times!

We assume in these cases that in case of tie, Strassel would win.

B.

Profit:

Case 1 = 0.309 * (155000-125000) = $9,259

Case 2 = 0.605 * (155000-135000) = $12,098

Case 3 = 1 * (155000-145000) = $10,000

Therefore recommended bid is $135,000 and the estimated profit is $12,098

Worksheet is given below. Simulated by random number up to 1000 rows. I have only given first few rows due to space constraint

random1 random2 bid1 bid2 win1 win2 win simulated wins total wins percent
1 0.355602 0.288527 4 3 1 1 1 583 554 307 31%
2 0.514042 0.547169 5 5 1 1 1
3 0.276026 0.727693 3 7 1 0 0
4 0.040496 0.526015 0 5 1 1 1
5 0.540871 0.82348 5 8 1 0 0
6 0.315083 0.081181 3 1 1 1 1 win profit
1 100000 7 0.859022 0.277518 9 3 0 1 0 25000 20000 45000 0.555556 0.308642 9259.259
2 105000 8 0.132149 0.08512 1 1 1 1 1 35000 10000 45000 0.777778 0.604938 12098.77
3 110000 9 0.632255 0.978488 6 10 0 0 0 45000 0 45000 1 1 10000
4 115000 10 0.915841 0.724907 9 7 0 0 0
5 120000 11 0.823775 0.174383 8 2 0 1 0
6 125000 12 0.083028 0.344359 1 3 1 1 1
7 130000 13 0.712683 0.295336 7 3 0 1 0
8 135000 14 0.067634 0.295011 1 3 1 1 1
9 140000 15 0.822325 0.095052 8 1 0 1 0
10 145000 16 0.733952 0.228798 7 2 0 1 0
17 0.117918 0.401072 1 4 1 1 1
18 0.334256 0.797895 3 8 1 0 0
19 0.163273 0.346425 2 3 1 1 1
20 0.697902 0.440001 7 4 0 1 0
21 0.462317 0.697812 5 7 1 0 0
22 0.858809 0.439994 9 4 0 1 0
23 0.646317 0.952174 6 10 0 0 0
24 0.813409 0.518206 8 5 0 1 0
25 0.219426 0.144305 2 1 1 1 1
26 0.692166 0.965413 7 10 0 0 0
27 0.363265 0.740157 4 7 1 0 0